You're driving down the highway late one night at 20m/s when a deer steps onto the road 57 m in front of you. Your reaction time before stepping on the brakes is 0.50s, and the maximum deceleration of your car is 10m/s2. What is the maximum speed you could have and still not hit the deer?
Answer #1

let the speed with which we can drive be u m/s

the safe distance we can have

d = 57 - 0.5 * u

niw using 2nd equation of motion

v^2 = u^2 - 2 a d

0 = u^2 - 2* 10* ( 57 - 0. 5* u)

u^2 + 10 u - 1140 = 0

solving for u

u = 29.132 m/s

=======

Comment in case any doubt, will reply for sure Goodluck